# Generalizing Scaling Laws for Dense and Sparse Large Language Models

### v2

@@ III Generalizing Scaling Laws
We observed that the previous work and their scaling laws are based on the number of nonzero or active parameters. Dense models use all the parameters, and hence in [3] $N$ was used as the total number of parameters in its scaling law. In [4] the nonzero parameters $N$ were considered in its scaling law 3. In [5] the authors modified the compute budget $C=6ND$ from [2] to $C=6N_{a}D$, where $N_{a}$ is the number of active parameters. These scaling laws do not cover the dense scaling law when the sparsity is 0.
These observations and limitations motivate us to generalize the scaling laws of [3, 4, 5] into a single representation that can effectively capture the experimental behavior and features observed in these works. We want to make sure that the new scaling law works for both dense and sparse models. As we increase model size, data size, and compute, we want to get better performance [2, 3, 4, 5]. Similarly, as we increase the sparsity, we want to obtain performance improvement [4, 5]. Furthermore, for $0\%$ sparsity or dense models, we need to make sure that our scaling law behaves the same as the original scaling law in Equation 2.
For simplicity, we use the term active parameters $N$ for active/nonzero parameters in the MoE and pruned models. For the dense models, the total number of active parameters is the total number of parameters. We use Equation 2 as the base to extend it to a generalized scaling law. We define the sparsity as follows:
@@
Where $0\leq S<1$ because the active parameters in a model cannot be zero. When $S=0$, it means that the total parameters is the same as the active parameters. Therefore, $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the approach in [4] we introduce an upper-bound constant sparsity factor $c$, which is caused by the sparsity; but this factor, which contributes to the loss, depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$. For sparsity $S=0$ or dense models, this term is the same as the original parameter-dependent term in Equation 2.
We have noticed that different approaches adjusted the entropy $e$ in different ways, as shown in Tables II, IV, and V. The value of $e$ is impacted by the sparsity. We adjust the value of $e$ by multiplying it with $(1-S)^{\gamma}$, where $\gamma$ is a constant. Accumulating these changes, we propose the following scaling law:
@@
This equation quantifies the impact of sparsity on loss. Notice that we did not change the values of the data-dependent term in Equation 2, because the work in [4, 5] has shown experimentally that sparsity has no effect on this term.
The fitted coefficient values in our proposed Equation 7 are listed in Table VI. Since we have used the empirical scaling law in [3] as the base, we simply apply the coefficient values in Table II to Equation 7. For the sparsity factor $c$, we use the dataset for the largest model with the largest sparsity in [4] to fit Equation 7 to obtain the value for the sparsity factor $c$. The value of $\gamma$ was empirically obtained.
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $1.69$ $406.4$ $410.7$ $93.45$ $0.34$ $0.28$ $10^{-2}$
TABLE VI: Fitted parameter values for Equation 7
Notice that Equation 7 is similar to Equation 4; however, the coefficients and the parameter-dependent term are different. If the sparsity is 0, Equation 4 is not the same as Equation 2, as shown in Figure 2.
Lemma: If the sparsity $S$ is 0, Equation 7 is the same as Equation 2.
@@ IV-A Performance Results
In this section we compare the performance between our proposed scaling law and the previous scaling laws for dense and sparse models. We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [3, 4, 5]. Hoffman et al. [3] provided a list of nine models with number of parameters and training tokens used in their experiments. Franta et al.  [4] provided four different numbers of nonzero parameters, three different numbers of training tokens, and four different sparsity levels with a total of $48$ experimental runs to collect their experimental results. Abnar et al.  [5] used five different numbers of compute budgets, seven different sparsity levels, and a range of model sizes. Using the formula $C=6ND$, we calculated their number of training tokens and listed the ranges in Table VII. We use these experimental datasets to generate plots for their respective scaling laws, and we use the same datasets to generate plots using our proposed scaling law in order to do a head-to-head comparison.
@@
Hoffman et al. [3] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
Frantar et al. [4] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
Abnar et al. [5] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
TABLE VII: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
When we compare our proposed scaling law in Equation 7 with the dense scaling law [3] in Equation 2, as we discussed in the lemma in the preceding section, if the sparsity is 0, Equation 7 is the same as Equation 2. Figure 4 shows the loss vs compute budget using the dataset in [3]. On the $x$-axis is compute. which depends on parameters $N$ and tokens $D$ and hence captures the impact of both ($C=6ND$). On the $y$-axis is the calculated loss. We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models.
Figure 4: Comparison between Hoffman’s scaling law in Equation 2 and the proposed scaling law in Equation 7
We compare the proposed scaling law in Equation 7 with the scaling law [4] in Equation 3. We use the 48 models to conduct the experiments in [4] for performance comparison. Figures 5 (a) and 5 (b) show that our proposed scaling law replicates the behavior of the scaling law in [4]. We observe that with the increase of the sparsity the loss decreases. However, there is a difference between the loss values for both scaling laws which is similar to Figure 2. This is because the models used for the experiments in [4] are relatively much smaller (the maximum number of parameters just 85M) than other works, and the adjusted coefficient values in the proposed scaling law are for larger models, as shown in Figure 4 where the loss decreases with increasing the model size and compute budget. Hence we see a relatively larger loss for the much smaller datasets.
(a) Scaling behavior for Equation 3 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 5: Comparison between (a) the scaling law in Equation 3 and (b) the proposed scaling law in Equation 7
We also compare our proposed scaling law in Equation 7 with the scaling law [5] in Equation 5. For performance comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models the authors used to conduct the experiments in [5]. We have drawn compute budget vs loss plots in Figures 6(a) and 6(b) to compare Equations 5 and 7. We observe that with the increase of the sparsity and compute, the loss decreases; and both figures show a similar pattern. Increasing the compute budget means an increase in the active parameter size ($N_{a}$) and number of tokens ($D$). This indicates that the proposed scaling law shows scaling behavior similar to that in [5].
(a) Scaling behavior for Equation 5 with varying compute budget
(b) Scaling behavior for Equation 7 with varying compute budget
Figure 6: Comparison between (a) the scaling law in Equation 5 and (b) the proposed scaling law in Equation 7
IV-B Hyperparameter Optimization for Better Coefficients
Most of the previous scaling laws discussed in this paper used grid search to find the best values for the coefficients. While grid search is a commonly used approach, it is still inefficient. As an alternative, we have used ytopt [36, 37, 38] for hyperparameter optimization. ytopt is a Bayesian optimization–based autotuner that builds a surrogate model to explore the promising regions to find the optimal values with fewer evaluations. In our previous work [39], ytopt outperformed grid search, random search, a genetic algorithm, and XGBoost in accuracy and tuning time. Unlike traditional autotuners, ytopt leverages libEnsemble’s [40] asynchronous manager-worker framework to run multiple evaluations in parallel, reducing the one-by-one autotuning time.
To use ytopt for the optimization process for finding the best coefficients for the previous scaling laws, we use the best coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient. We carefully define the range for each coefficient for Equations 2, 3, and 5 and then use ytopt to run the optimization process. The generated hyperparameter optimization results in Figure 7(a), (b), and (c) illustrate that ytopt identifies the best coefficient configuration (combination of coefficients) over time (blue) with the baseline (red) for the default values and that using ytopt results in lower loss values than the coefficient-based loss values (red, baseline) using grid search.
@@ IV-B Hyperparameter Optimization for Better Coefficients
Figure 7: Loss calculated by the coefficients generated by ytopt and grid search. The red lines are the loss using the best coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
@@ IV-C Discussion
In summary, we observe that our proposed scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs. In Figure 8 the proposed scaling law shows no spike for higher sparsity with the fixed compute budget 1e20 as we observed for Equation 3 in Figure 1. This indicates that the proposed scaling law captures the scaling behaviors for higher sparsity. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with the given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt [36, 37].
In Figure 5 we observe that the range of loss values is different for Equations 3 and 7. This difference is mainly because the experiments done for Equation 3 use much smaller model sizes with at most 85M. For the scaling law in Equation 5 we can see that the results are almost identical. These results further confirm that our proposed scaling law captures the scaling behaviors of the MoE models.
Figure 8: IsoFLOP plot of the proposed scaling law in Equation 7 for different sparsities
In this work we have generalized the scaling behaviors of dense and sparse models. Our generalized scaling law extends and covers several existing scaling laws. Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger models. There are other ways [41, 42] to introduce sparsification in a model. Different training approaches [30, 32] may yield different scaling behavior as well. Further work can be done to incorporate these scaling laws into our generalized scaling law.
@@ V Conclusions
Current LLMs result in higher accuracies and exhibit few-shot learning and even human-like emergent abilities on a wide range of language tasks. However, these models are costly to train. As LLMs get larger and more expensive, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. But while architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated, pruned, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures. For a given compute budget, the proposed scaling law can be used to estimate the best model hyperparameters with a given sparsity or to identify the optimal sparsity for the given model hyperparameters by using ytopt.
Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [30, 31, 32] but also take into account inference [43], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting. Since the behavior of scaling laws varies based on training and inference approaches, these approaches can be incorporated in the generalized scaling law as well. Unifying different scaling laws into one single representation lays the foundation of architecture-agnostic representation of the scaling behavior of LLMs. As LLM models continue to diversify, this generalized approach will allow us to anticipate the return based on their allocated compute resources.
@@ Acknowledgments
This work was supported DOE ASCR RAPIDS2 and OASIS. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.

### v3

@@ III Generalizing Scaling Laws
We observed that regardless of the dense or sparse architecture, all the previous works mentioned so far use the number of nonzero or active parameters to measure the total compute. Hoffman et al. [20] used $C=6ND$ where $N$ is the total number of parameters and for dense models this is the number of nonzero/active parameters. Abnar et al. [1] used $C=6N_{a}D$ where $N_{a}$ is the number of active parameters. Frantar et al. [16] used $C=6ND\cdot c_{mul}(S)$ where the authors used dense Hoffman’s compute formula and multiplied it by a sparsity-dependent factor $c_{mul}(S)$. We have also observed in the preceding section that the mentioned sparse scaling laws had their limitations and do not replicate the dense scaling law behavior when the sparsity is $0\%$. In Section I we have also discussed that due to the architecture-specific nature of existing scaling laws where some scaling laws take total parameters into account for their scaling law and some use the number of nonzero/active parameters, cross-architecture comparison and budget planning become challenging. On top of that, different scaling laws use different scaling law coefficients that need to be tuned carefully to fit the early experimental results for better prediction. We think that the scaling behavior, regardless of the architecture (dense or sparse), is dependent on the number of active/non-zero parameters and propose a generalized scaling law that can address the earlier-mentioned issues.
This proposed scaling law should 1) be able to describe the scaling behavior for both dense and sparse models; 2) achieve similar predictions and capture the scaling behavior observed in the experiments of [23, 20, 16, 1] as increasing model size, data size, and compute; and 3) capture the behavior for sparse models as observed in [16, 1] for increasing the sparsity. Furthermore, for $0\%$ sparsity or dense models, the scaling law should behave exactly as the Hoffman scaling law.
Figure 5: Scaling behavior and prediction comparison between Abnar scaling law and our proposed scaling law.
For simplicity, we use the term “active parameters” ($N$) for active/nonzero parameters in the MoE and pruned models. For the dense models, the total number of active parameters is the total number of parameters. We use Hoffman scaling law (Equation 2) as the base to propose a generalized scaling law. We define the sparsity as follows:
@@
Where $0\leq S<1$ (The number of active parameters in a model cannot be zero and hence, $S\not=1$). $\frac{N}{1-S}$ is the total number of parameters. In the parameter-dependent second term of Hoffman scaling law in Equation 2, we replace $N$ with $\frac{N}{1-S}$; the parameter-dependent term thus becomes $\frac{a}{\left({\frac{N}{1-S}}\right)^{\alpha}}$, which is simplified as $\frac{a(1-S)^{\alpha}}{N^{\alpha}}$. Similar to the Frantar et al. approach in [16] we introduce an upper-bound constant sparsity factor $c$; but unlike Frantar scaling law, this factor depends on the sparsity $S$. So we define the parameter-dependent term as $\left(a(1-S)^{\alpha}+c\cdot S\right)\frac{1}{N^{\alpha}}$.
While $e$ corresponds to irreducible loss, empirical fits over $N$ and $D$ treat the constant as an effective offset. For sparse models, this offset can vary systematically with sparsity. We model this dependence with minimal multiplicative factor $(1-S)^{\gamma}$, preserving the dense limit and power law behavior of scaling laws. Accumulating these changes, we propose Equation 7 as our proposed scaling law.
@@
This equation quantifies the impact of sparsity on loss. Notice that we did not change the values of the data-dependent term ($\frac{b}{D^{\beta}}$) in Equation 2 (Hoffman scaling law), because the work in [16, 1] has shown experimentally that sparsity has no effect on this term.
Figure 6: IsoFLOP plot of Abnar scaling law (left) and our proposed scaling law (right) for a compute budget of $1e20$ FLOPs.
Lemma: If sparsity $S$ is 0, Equation 7 is the Hoffman scaling law in Equation 2.
@@
Figure 7: IsoFLOP plot of Abnar scaling law (left) and our proposed scaling law (right) for a compute budget $3.29e24$ FLOPs of DeepSeek-V3 training.
@@ IV-A Performance Results
In this section we compare the performance of our proposed scaling law and the previous architecture specific scaling laws for dense and sparse models. We have collected the experimental datasets (number of parameters, number of tokens, and sparsity) used in [20, 16, 1] shown in Table VI. Hoffman et al. [20] provided a list of nine models with number of parameters and training tokens used in their experiments. Frantar et al.  [16] provided four different numbers of nonzero parameters, three different numbers of training tokens, and four different sparsity levels with a total of $48$ experimental runs to collect their experimental results. Abnar et al.  [1] used five different numbers of compute budgets, seven different sparsity levels, and a range of model sizes. Then we use these experimental datasets to generate plots for their respective scaling laws and use the same datasets to generate plots using our proposed scaling law in Equation 7 in order to compare them.
@@
Hoffman [20] 400M10T400M{-}10T 8B216.2B8B{-}216.2B - 99
Frantar [16] 1.3M85M1.3M{-}85M 16B65B16B{-}65B $0,50,75,87.5$ $48$
Abnar [1] 329M21.2B329M{-}21.2B 15B128B15B{-}128B $0,25,50,75,90,95,98$ $35$
TABLE VI: Parameters, tokens, sparsity, and number of test models collected from the corresponding works.
Along with drawing the plots, we measure the prediction Mean Square Error (MSE) between our proposed scaling law and the respective scaling law for better visualization of prediction of our proposed scaling law.
@@
Figure 8: Loss calculated by the scaling law coefficients generated by ytopt and grid search. The red lines are the loss using the best scaling law coefficients provided in the respective work, and the blue dots are the loss tuned for different evaluations over time using ytopt.
First, to compare our proposed scaling law in Equation 7 with Hoffman scaling law in Equation 2 [20], we refer back to the Lemma in section III where we have shown that if the sparsity $S=0$, Equation 7 converts into Equation 2. Table VII shows the fitted coefficient values in Equation 7 for the dense models from [20]. Figure 3 shows the loss vs compute budget using the dataset in [20]. On the $x$-axis is compute ($C=6ND$). which depends on model size $N$ and tokens $D$ and hence captures the impact of both. On the $y$-axis is the calculated loss using both scaling laws. We see an identical plot, indicating that our proposed scaling law behaves the same as the original scaling law for dense models capturing the dense scaling behavior where increasing compute results in decreasing pretraining loss.
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $1.69$ $406.4$ $410.7$ 0 $0.34$ $0.28$ $0.01$
TABLE VII: Fitted scaling law coefficient values of our proposed scaling law for dense models.
To compare our proposed scaling law with Frantar scaling law, we use the data (sparsity, training steps, number of non-zero parameters and training tokens) collected from the 48 models that were used to conduct the experiments in [16]. Figure 4 shows that our proposed scaling law captures the almost similar scaling behavior as Frantar scaling law for sparse models as sparsity and number of of nonzero parameters keep increasing. The loss prediction is very similar with average $MSE=0.00078$. Table VIII shows the fitted parameter values for our proposed scaling law for models from Frantar scaling law experiments [16].
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $0.15$ $86.03$ $7.90$ $29.26$ $0.28$ $0.08$ $2.00$
TABLE VIII: Fitted parameter values of our proposed scaling law for Frantar-style sparse models.
We also compared our proposed scaling law with the Abnar scaling law in Equation 5 [1]. For this comparison we use the interpolated values of parameters, tokens, and the exact compute budget of the 35 models used to conduct the experiments in [1]. We have used interpolated values to generate the full dataset because even though the authors have shared a list of compute budget and sparsity levels, no such list was shared for model size and training tokens. So we used the plots shared in [1] and generated the entire dataset. In Figure 5 we have drawn compute budget (FLOPs) vs. loss plots to compare the behavior of both scaling laws. We observe that with the increase of the sparsity and compute, the loss decreases, and both scaling laws show a similar pattern. The average prediction MSE between Abnar scaling law and our proposed scaling law is $0.0005$ which is not significant. We present our fitted scaling law coefficient values in Table IX.
Coefficients $e$ $a$ $b$ $c$ $\alpha$ $\beta$ $\gamma$
Values $0.57$ $8.26$ $6324.82$ $3.57$ $0.08$ $0.40$ $1.19$
TABLE IX: Fitted parameter values of our proposed scaling law for MoE-based models.
Abnar et al. [1] also conducted IsoFLOP experiments with the given compute budget of $1e20$ FLOPs for the MoE-based models. In Figure 6 we have shown the IsoFLOP loss prediction comparison between Abnar scaling law and our proposed scaling law, where each line connects models with different sparsity and model sizes trained under the same amount of compute budget $1e20$ with active parameters as x-axis and training loss as y-axis. In [1] the authors intentionally cut the plot for active parameters at $4B$ parameters even though the original experiments had active parameters around $26B$. For comparison purpose, we mark that $4B$ parameter spot and generate the entire plot in Figure 6. For IsoFLOP experiments [1], the authors showed that for fixed sparsity, as we increse number of active parameters, the pretraining loss keeps decreasing until it reaches a sweet spot. Beyond that sweet spot, loss starts increasing again. As seen from Figure 6, our proposed scaling law can effectively capture this behavior.
To demonstrate the effectiveness of our scaling law for larger models, we use DeepSeek-V3 [11] which is a MoE-based model with 671B parameters and 37B of which are activated for each token as an example. The model was trained with $14.8T$ tokens. Using the compute formula $C=6ND$ where $N$ is the number of active parameters and $D$ is the number of training tokens, we measure the total compute, $3.29e24$ used to train DeepSeek-V3. We generate an interpolated IsoFLOP dataset using this compute budget where we restricted the number of total parameters to $671B$ parameters. Figure 7 shows that our proposed scaling law exhibits a very similar pattern to Abnar scaling law for very large models as well.
IV-B Hyperparameter Optimization for Better Scaling Law Coefficients
Most of the previous scaling laws discussed in this paper used grid search to find the best values for the scaling law coefficients. While grid search is a commonly used approach, it is still inefficient. As an alternative, we have used ytopt [39, 41, 42] for scaling law coefficient optimization. ytopt is a Bayesian optimization–based autotuner that builds a surrogate model to explore the promising regions to find the optimal values with fewer evaluations. In our previous work [40], ytopt outperformed grid search, random search, a genetic algorithm, and XGBoost in accuracy and tuning time. Unlike traditional autotuners, ytopt leverages libEnsemble’s [21] asynchronous manager-worker framework to run multiple evaluations in parallel, reducing the one-by-one autotuning time.
To use ytopt for the optimization process for finding the best scaling law coefficients for the previous scaling laws, we use the best scaling law coefficients for the scaling laws in Tables II, IV, and V as the baseline and the default values to define the proper search space for each coefficient. We carefully define the range for each coefficient for Equations 2, 3, and 5 and then use ytopt to run the optimization process. The generated hyperparameter optimization results in Figure 8(a), (b), and (c) illustrate that ytopt achieves better loss predictions than grid search. Since ytopt is a Bayesian optimization-based approach, it is much faster than a grid search-based approach and is overall a better alternative.
@@ IV-C Discussion
In summary, we observe that our proposed generalized scaling law Equation 7 effectively captures the scaling behavior of Equations 2, 3, and 5 for dense and sparse LLMs. The effectiveness is further justified by the small MSE between the existing scaling law and out proposed scaling law predictions. In Figure 6 and 7 we have seen that our proposed scaling law can capture the IsoFLOP behavior of MoE models as well. So we can conclude that for a given compute budget, our proposed scaling law can be used to estimate the best model hyperparameters (Model size, Tokens and Compute) for a given sparsity or to identify the optimal sparsity for the given model hyperparameters.
In this work we have generalized the scaling behaviors of dense and sparse models. Our generalized scaling law extends and covers several existing scaling laws. Its limitations are related to the sparsity factor $c$ and the coefficient $\gamma$, which need to be further optimized for larger LLM models.
@@ V Conclusions
As LLMs get larger and more expensive in training, it is crucial for researchers and developers to estimate the expected performance beforehand. Scaling laws thus are critical. While architecture-specific scaling laws give a fine-grained estimation, as more and more new architecture and training methods are being developed, using different scaling laws for each variant becomes impractical. Our generalized scaling law covers densely activated models, pruned models, and MoE models. Empirical results show that our generalized approach can effectively capture the scaling behavior of the architectures.
Because LLMs inference can be memory-bound and compute-intensive, future work will extend the generalized scaling law in training to not only incorporate other scaling laws [24, 4, 5, 38, 35] but also take into account inference [32], especially recent agentic AI workloads for answering complex questions through reasoning using chain-of-thought prompting. Since the behavior of scaling laws varies based on training and inference approaches, these approaches can be incorporated in the generalized scaling law as well. Unifying different scaling laws into one single representation lays the foundation of architecture-agnostic representation of the scaling behavior of LLMs.
@@ Acknowledgments
This work was supported by DOE ASCR SciDAC RAPIDS and OASIS. We acknowledge the Argonne Leadership Computing Facility for use of Polaris under the project EE-ECP. This material is based upon work supported by the U.S. Department of Energy, Office of Science, under contract number DE-AC02-06CH11357.